# Rethinking Visual Layer Selection in Multimodal LLMs

### v1

@@ Abstract
Multimodal large language models (MLLMs) have achieved impressive performance across a wide range of tasks, typically using CLIP-ViT as their visual encoder due to its strong text-image alignment capabilities. While prior studies suggest that different CLIP-ViT layers capture different types of information—shallower layers focusing on fine visual details and deeper layers aligning more closely with textual semantics—most MLLMs still select visual features based on empirical heuristics rather than systematic analysis. In this work, we propose a Layer-wise Representation Similarity approach to group CLIP-ViT layers with similar behaviors into shallow, middle, and deep categories and assess their impact on MLLM performance. Building on this foundation, we revisit the visual layer selection problem in MLLMs at scale, training LLaVA-style models ranging from 1.4B to 7B parameters. Through extensive experiments across 10 datasets and 4 tasks, we find that: (1) deep layers are essential for OCR tasks; (2) shallow and middle layers substantially outperform deep layers on reasoning tasks involving counting, positioning, and object localization; (3) a lightweight fusion of features across shallow, middle, and deep layers consistently outperforms specialized fusion baselines and single-layer selections, achieving gains on 9 out of 10 datasets. Our work offers the first principled study of visual layer selection in MLLMs, laying the groundwork for deeper investigations into visual representation learning for MLLMs.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) [38, 39, 40, 48, 3] have catalyzed the development of Multimodal Large Language Models (MLLMs) [12, 3, 5], which aim to unify visual understanding with language reasoning – a capability foundational to applications ranging from robotic navigation to medical diagnosis [12, 2, 6, 3, 46]. While the integration of advanced LLMs has significantly enhanced the linguistic prowess of MLLMs, the vision component remains a critical yet underexplored bottleneck. Specifically, the selection of optimal visual features from hierarchical layers of Vision Transformers (ViTs) is often empirically driven [49, 16, 47] , lacking systematic justification.
Current MLLMs predominantly adopt a deep-layer bias: models like LLaVA [27, 28, 26] extract visual features from the penultimate layer of CLIP-ViT [40], while InternVL [6] variably uses the final or fourth-to-last layer, e.g. versions like InternVL 6B v1.2 and v1.5 use the final layer, while other variants within its series opt for the fourth-to-last layer as their visual representation. These choices, however, are rarely grounded in rigorous analysis. Prior works [10, 49, 16, 47, 4, 5] have demonstrated that ViT layers encode distinct semantic hierarchies (e.g., edges in shallow layers vs. object parts in deeper layers), yet their impact on MLLM performance remains unclear. This raises a pivotal question: Does the conventional reliance on deep layers truly maximize MLLM capabilities, or are we overlooking untapped potential in shallow/middle layers?
Although integrating features from multiple hierarchical levels has been extensively studied in computer vision [13, 25], its systematic exploration in the context of MLLMs is still lacking. While works like CLIP-InterpreT [10] reveal that different attention heads within ViT layers attend to distinct properties, there has been limited investigation into how such internal variations should inform the selection or fusion of visual features in multimodal models.
@@
To address this gap, we propose Layer-wise Representation Similarity (LRS) to quantify the cosine similarity of hidden states across CLIP-ViT layers. Our analysis reveals three distinct visual representation spaces: shallow (layer 1-12), middle (layer 13-20), and deep (layer 21-24), as illustrated in Fig. 1.
Building upon this foundation, we first re-evaluate the effectiveness of different deep layers. Although the penultimate layer does not always yield the absolute best performance, we find that it consistently outperforms other deep layers across a range of LLM scales (1.4B, 2.7B, and 7B parameters). This superiority stems from its ability to preserve more visual details while maintaining strong text alignment, validating its widespread use. However, we also observe that as LLM size increases, gains diminish when visual quality becomes the bottleneck—particularly in hallucination-prone tasks. This highlights an important limitation: the final-layer degradation caused by CLIP’s contrastive training cannot be compensated for simply by scaling up the LLM. Thus, among deep layers, the penultimate layer remains the best default choice.
Having established the penultimate layer’s dominance within the deep space, we turn to an even more critical question: could earlier layers offer complementary strengths? Surprisingly, our experiments reveal that shallow and middle layers outperform deep layers on approximately one-third of the MME [9] sub-tasks, particularly those involving counting and positioning. Consistent patterns are observed in MMVet [51]. As shown in Tab. 1, representations from the layer 18 surpass those from the penultimate layer by 3% on CVBench [46]. Fig. 2 shows that, although shallow layers have the worst overall performance compared to the middle and deep layers, it still outperforms them in a notable number of sub-tasks. To validate the robustness of these findings, we evaluate performance across three data scales (665k, 737k, and 1M samples). Despite some task-specific variations, the core conclusion remains: shallow and middle layers harbor untapped potential.
Motivated by these insights, we propose a minimalist fusion strategy that integrates features from shallow, middle, and deep layers. Our design employs only a single linear layer, incurring minimal computational overhead while substantially improving visual representations. This simple yet effective approach demonstrates that lightweight fusion strategies can systematically unlock performance gains, offering a principled alternative to ad-hoc layer selection in future MLLM research. Although previous studies [49, 12, 7, 4] have explored hierarchical feature fusion or LLM-aligned feature selection, they often rely on empirical heuristics without a dedicated analysis of the intrinsic layer properties themselves. In contrast, our work is the first to systematically analyze information variation across ViT layers and how it aligns with LLMs under controlled experimental setups.
The key contributions of this paper can be summarized as follows:
To analyze visual layer relationships in CLIP-ViT, we propose LRS, identifying three distinct groups: shallow (1–12 layers), middle (13–20 layers), and deep (21–24 layers). (Sec. 3.3)
- (2)
- (3)
@@ 2 Related Work
2.1 Multimodal LLMs
The success of large language models (LLMs) [39, 48, 8, 3] and pretrained visual models [45, 40, 52] in their respective fields have laid a solid foundation for the emergence of multimodal LLMs. For instance, LLMs such as the LLaMA [8] series demonstrate powerful natural language processing capabilities, while CLIP-ViT [40] align the visual and textual modalities effectively through large-scale image-text contrastive learning.
Thanks to the contributions of numerous open-source language models, the parameters of LLMs [8] have expanded from 7B to 110B, significantly boosting multimodal performance. Additionally, some studies have focused on improving the visual capabilities of these models, with methods like dynamic splitting being employed to better understand high-resolution images.
Inspired by recent advances in computer vision, such as DenseNet [13] and FPN [25], some studies have explored integrating shallow and deep features within the ViT of multimodal language models. However, these approaches implement feature integration without a detailed analysis of layer-wise differences in multimodal tasks.
2.2 Vision Encoder in MLLMs
@@
Serving as the “eyes” of Multimodal Large Language Models (MLLMs), the vision encoder sets the upper limit of the model’s perceptual capabilities. If certain information is absent from the representation, theoretically, the model cannot correctly answer questions that rely on this information as prior knowledge. CLIP, through image-text contrastive learning effectively aligns visual representation with text space and is widely adopted as the visual encoder in models such as LLaVA [27, 26], Qwen-VL [3], Flamingo [2], and BLIP [23]. Other foundational vision models, such as DINOv2 [37], SigLIP [52], ConvNeXT [32], are also utilized to build MLLMs. In this paper, we selected the widely used CLIP-ViT model as the focus of our layer-wise analysis.
2.3 Visual Fusion Methods
Visual feature fusion methods can be distinguished by whether visual tokens interact directly with text tokens. One approach [49, 47] merges hierarchical visual features through element-wise or channel-wise fusion before feeding them into the connector module, rather than directly using a single-layer ViT.
Another approach [2, 12, 5] uses a cross-attention mechanism to enable interaction between visual and textual representations, resulting in stronger text perception and comprehension capabilities.
However, understanding the differences in visual representations generated by different encoders remains a key research question in multi-modal large models. This study aims to provide a comprehensive analysis of layer-wise visual representations in MLLMs, offering valuable insights for the development of future visual fusion methods.
3 Model, Benchmark, and Analysis Method
3.1 Base Model
MLLMs have various structural designs, and our analysis of the role of visual layers is based on an LLaVA-like model. In this section, we will provide a detailed illustration of the structure of LLaVA-like models and the details of our implementation process.
LLaVA-like MLLM Structure
As one of the leading architectures in modern MLLMs, LLaVA consists of a CLIP-based visual encoder, a projection layer, and a pre-trained LLM. The encoder extracts image features and maps them to the LLM’s token space, enabling seamless vision-language interaction.
To explore the role of different visual layers in MLLMs, an ideal approach is to input each layer into a full MLLM and evaluate its performance on multimodal benchmarks. Given our focus on the visual side, a more efficient alternative is to use an LLaVA-like structure with a smaller LLM for layer-wise experiments.
In this equation, represents the initial hidden state of the model, which is derived from the input image through an embedding function . The input denotes a image with height , width , and channels.
Each hidden state is recursively computed using a TransformerBlock applied to , where and is the number of layers. Each TransformerBlock consists of multi-head self-attention (MSA) and feed-forward (FFN) layers, progressively refining feature representations. Our study focuses on analyzing these hidden states, termed visual representations, in the context of multimodal tasks.
Here, denotes the visual embedding, and denotes the text embedding. Each hidden state captures a different level of semantic information, with deeper layers generally encoding more abstract, high-level features. In our layer-wise experiment, is the input of the connector in MLLMs. Each visual representation space demonstrates unique characteristics across various tasks. In Sec 4, we provide a comprehensive analysis of our findings.
Implementation Details
We employ CLIP ViT-L/14 (336px) [40] as the visual encoder and 1.4B MobileLLaMA as the language model for efficiency analysis, with a one-layer MLP serving as the connector. Training follows a two-phase strategy aligned with LLaVA [27]. AdamW optimizer with a cosine annealing scheduler is used, with learning rates of 1e-3 (phase one) and 2e-5 (phase two), and batch sizes of 256 and 128. Training on four NVIDIA A100 80GB GPUs takes 2 hours for phase one and 8 hours for phase two. We adopt the LLaVA 1.5 [27] dataset, comprising 558K image-caption pairs for pre-training and 665K conversational instances for instruction tuning. Unless explicitly noted, the experimental setup remains the same.
@@ 3.1 Base Model
3.2 Evaluation Benchmarks
To comprehensively explore and evaluate various visual representations, we classified the benchmarks into four categories following previous work [46]: (1) General tasks, (2) OCR (Optical Character Recognition) tasks, (3) Vision-centric tasks, and (4) Hallucination tasks.
The General tasks category includes MME [9] (yes/no questions on attributes like existence and color), MMBench [30] (multiple-choice across diverse aspects), SEEDBench [22] (spatial and temporal reasoning), and GQA [15] (complex real-world VQA). The OCR category features TextVQA [44] and OCRBench [31], both assessing OCR capabilities. The Vision-centric category includes CVBench [46] (evaluating spatial relations and depth), RealWorldQA (real-world QA), MMVet [51] (general multimodal assessment), and RefCOCO [50] (visual grounding). Finally, the Hallucination category includes POPE [24], which evaluates object hallucination.
3.3 Layer-wise Representation Similarity
To examine the behavioral patterns of different visual layers, we propose an analysis method named Layer-wise Representation Similarity (LRS) to analyze the relationships between various visual layers.
Specifically, we compute the cosine similarity between hidden states of ViT layers, where values closer to indicate higher similarity. Fig. 1(a) shows the average cosine similarity of different visual layers. The formalized division process is detailed in the following equation:
We define the hidden state matrix of CLIP-ViT as:
From Fig.1(a), we can identify three distinct representation spaces. As illustrated in Fig.1(b), shallow layers contribute minimally to OCR tasks, performance improves rapidly in the middle layers, and reaches its peak in the deep layers. This pattern suggests that visual layers within the same partition exhibit similar behaviors.
@@ 4 Experiment: Layer-wise Exploration
In this section, we address Q1: Can shallow and middle visual representations outperform deep one? We conduct a layer-wise analysis and evaluate across 10 benchmarks to compare different representations.
4.1 Settings
Layer-wise Analysis
To fully explore each layer’s potential, we input representations from each layer of the CLIP-ViT model into the connector. We conduct a systematic evaluation of visual representations across multimodal benchmarks. Specifically, we take through from Sec. 3 as inputs, followed by two-stage training. This setup helps assess the upper bound of visual representations in multimodal tasks.
4.2 Exp-I: Effect of Shallow and Middle Layers
Potential of Shallow and Middle layers
We first conduct an investigation into the shallow and middle representation spaces. The performance of these two spaces as shown in Tab. 1, several key insights emerge from the results:
The middle layer has the potential to perform best: Although the middle layer’s information has not been fully processed, it still achieves the best performance on one-third of the benchmarks. Specifically, compared to the penultimate layer, layer 14 achieves a 20-point higher score on MME-Cognitive, layer 18 outperforms by 3% on CVBench, layer 17 surpasses by 1.4% on RefCOCO, and layer 20 exceeds by 0.2% on GQA. This demonstrates the potential of middle layers.
The middle layers generally perform better on vision-centric tasks: Fig. 3(a) illustrates the performance of different layers across subtasks in the MME dataset, showing that position and existence tasks benefit more from middle layer representations. As depicted in Fig. 3(b), the penultimate layer achieves top performance in only three out of eleven subtasks, whereas the shallow and middle layers yield optimal results in seven. Similarly, in Fig. 3(c), one-third of the best performing results, such as those in spatial relations, physical relations, and cross fine-grained perception originate from the shallow and middle layers. A comparable trend is observed in SEEDBench (Fig. 3(d)), where middle layers produce optimal results in nearly half of the subtasks, including Instance Attribute, Instance Location, Instance Interaction, and Text Recognition.
Hallucination issues can be effectively handled at the middle layers: As shown in Tab. 1, POPE results indicate that hallucination issues are most prominent in the shallow representation space, with minimal variation between the middle and deep layers. Notably, in the middle representation space, half of the layers outperform the penultimate layer on the POPE. This is because, comparatively, the challenge of the task lies more in perception than comprehension. In Sec. 5.3, we provide a detailed analysis showing that further experiments with larger LLMs consistently support this finding.
4.3 Exp-II: Effect of Deep Layer
Which layer performs the best in deep space?
A widely accepted but rarely questioned assumption is which layer of visual representation should be used as input for MLLMs. Our findings reveal that utilizing the final layer visual representation as the input is not the optimal choice and may not even be a close second:

### v2

# Multimodal Language Models See Better When They Look Shallower

@@ Abstract
Multimodal large language models (MLLMs) typically extract visual features from the final layers of a pretrained Vision Transformer (ViT). This widespread deep-layer bias, however, is largely driven by empirical convention rather than principled analysis. While prior studies suggest that different ViT layers capture different types of information—shallower layers focusing on fine visual details and deeper layers aligning more closely with textual semantics, the impact of this variation on MLLM performance remains underexplored. We present the first comprehensive study of visual layer selection for MLLMs, analyzing representation similarity across ViT layers to establish shallow, middle, and deep layer groupings. Through extensive evaluation of MLLMs (1.4B–7B parameters) across 10 benchmarks encompassing 60+ tasks, we find that while deep layers excel in semantic-rich tasks like OCR, shallow and middle layers significantly outperform them on fine-grained visual tasks including counting, positioning, and object localization. Building on these insights, we propose a lightweight feature fusion method that strategically incorporates shallower layers, achieving consistent improvements over both single-layer and specialized fusion baselines. Our work offers the first principled study of visual layer selection in MLLMs, showing that MLLMs can often see better when they look shallower.
Multimodal Language Models See Better When They Look Shallower
Haoran Chen1,2,3††thanks: Equal contribution., Junyan Lin2,311footnotemark: 1, Xinghao Chen2,3, Yue Fan4, Jianfeng Dong1, Xin Jin2,3, Hui Su5, Jinlan Fu6, Xiaoyu Shen2,3††thanks: Corresponding Author. Code available at https://github.com/EIT-NLP/VisualProbing-for-MLLM. 1Zhejiang Gongshang University 2Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative 3Institute of Digital Twin, Eastern Institute of Technology, Ningbo 4Genmo.ai, 5Meituan Inc., 6National University of Singapore haoranchr321@gmail.com, xyshen@eitech.edu.cn
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) extend the capabilities of traditional Large Language Models (LLMs) by enabling joint reasoning over both visual and textual inputs Hong et al. (2024); Bai et al. (2023); Chen et al. (2024a); Shen et al. (2024). Typically, these models integrate a pretrained Vision Transformer (ViT) to extract image features, which are then projected into the language embedding space of an LLM. This architecture enables unified multimodal understanding and powers a wide range of applications, including robotic navigation, medical diagnostics, and visual question answering Alayrac et al. (2022); Chen et al. (2024b); Bai et al. (2023); Tong et al. (2024a); Li et al. (2025).
While recent advancements have notably improved the language reasoning capabilities of MLLMs, the visual processing pipeline, particularly the selection of ViT layers for visual representation, remains insufficiently explored. In practice, MLLMs often default to using features from the deepest layers of ViT models. For instance, Qwen-VL Bai et al. (2025) and InternVL-6B v1.2/1.5 use the final layer of CLIP-ViT Radford et al. (2021), while other InternVL variants select the fourth-to-last layer Chen et al. (2024b). The LLaVA series Liu et al. (2023b, a, 2024a) relies on the penultimate layer. However, these choices are largely heuristics rather than systematic evaluation Yao et al. (2024); Jiang et al. (2023); Tong et al. (2024b); Cao et al. (2024); Li et al. (2026).
Previous work has shown that ViT layers encode a hierarchy of semantic information—from low-level edge detectors in shallow layers to abstract object representations in deeper layers Gandelsman et al. (2024); Yao et al. (2024); Tong et al. (2024b). Yet, how these layer-wise representations affect MLLM performance remains poorly understood. This paper addresses this gap by systematically investigating which ViT layers provide the most effective visual features for MLLMs.
@@
We begin by analyzing Layer-wise Representation Similarity (LRS) across CLIP-ViT’s hidden states using cosine similarity, revealing three semantically coherent layer groups: shallow (layers 1–12), middle (13–20), and deep (21–24) (Fig. 1). This categorization provides a foundation for structured layer selection and fusion.
Building upon this foundation, we first systematically assess the efficacy of different deep vision layers. Our analysis reveals that while the penultimate layer does not universally achieve peak performance in every scenario, it demonstrates consistent superiority across all evaluated model scales (1.4B, 2.7B, and 7B parameters). This advantage stems from the penultimate layer’s unique balance of preserving fine-grained visual details while maintaining strong alignment with textual representations. Notably, the performance gap between the penultimate layer and other deep layers widens as model scale increases. This suggests that simply using larger LLMs cannot compensate for suboptimal visual feature selection, underscoring the critical importance of visual layer choice in MLLMs.
Having established the penultimate layer’s strength among deep layers, we ask a more fundamental question: Can shallower ViT layers offer complementary or even superior information? Our analysis shows that shallow and middle layers outperform deep layers in approximately one-third of sub-tasks in the MME benchmark Fu et al. (2024) (Fig. 3), particularly in tasks involving fine-grained localization and counting. For instance, layer 18 outperforms the penultimate layer by 20% on position tasks (Fig.9). Similar trends are observed in MMVet Yu et al. (2023). Although shallow layers generally show lower average performance, they still excel on a significant subset of tasks (Fig.2). In contrast, deeper layers remain crucial for tasks with high-level semantic demands such as OCR. To assess robustness, we evaluate across three training data scales (665k, 737k, and 1M samples). Despite some fluctuations, our findings consistently demonstrate that shallow and middle layers carry underutilized yet valuable information.
Motivated by these insights, we propose a simple yet effective fusion strategy that combines visual features from shallow, middle, and deep layers. Our method uses a single linear projection layer, keeping computational overhead minimal while achieving substantial performance gains. This minimalist approach offers a principled alternative to existing ad-hoc layer selection and fusion methods. Unlike prior works Yao et al. (2024); Hong et al. (2024); Cao et al. (2024) that explore hierarchical feature fusion or LLM-aligned selection heuristically, our study provides the first systematic analysis of layer-wise information variation within ViTs, grounded in both intrinsic representation structure and downstream performance. Our key contributions are summarized as follows:
We identify three semantically coherent groups of ViT layers (shallow, middle, deep) based on representation similarity. We show that shallow and middle layers, which are often overlooked, can outperform the commonly used deep layers (Sec. 4).
Through extensive experiments across different data sizes and model scales, we confirm the generalization of our findings. Even as gains diminish with scaling, shallow and middle layers continue to exhibit unique strengths over deep layers in certain sub-tasks (Sec. 5).
We design a linear-layer-based fusion method that integrates features from all three layer groups. It outperforms both specialized fusion designs (e.g., DenseConnector Yao et al. (2024), MMFuser Cao et al. (2024)) and standard practices in current MLLMs (e.g., using only the penultimate layer) (Sec. 6).
@@ 2 Related Work
Visual Encoder in Multimodal LLMs
@@
Serving as the “eyes” of MLLMs, the vision encoder sets the upper bound of the model’s perceptual capabilities. CLIP, through image-text contrastive learning effectively aligns visual representation with text space and is widely adopted as the visual encoder in models such as LLaVA Liu et al. (2023b, a), Qwen-VL Wang et al. (2024), Flamingo Alayrac et al. (2022), and BLIP Li et al. (2023b). Other foundational vision models, such as DINOv2 Oquab et al. (2023), SigLIP Zhai et al. (2023), ConvNeXT Liu et al. (2022), are also utilized to build MLLMs. In this paper, we select the widely used CLIP-ViT model as the focus of our layer-wise analysis.
Visual Layer Selection
Recent studies have explored incorporating shallow visual features within the ViT of multimodal language models, such as DenseConnector Yao et al. (2024), MMFuser Cao et al. (2024). Lin et al. (2025) have further investigated internal fusion strategies by integrating multiple visual layers with language representations, highlighting the critical role of visual layer selection in effective multimodal integration.
Previous methods have largely relied on intuitive, heuristic-based strategies, such as evenly sampling layers. Although some approaches have explored the distinct characteristics of different ViT layers Gandelsman et al. (2024), the specific roles of layers at different depths in multimodal tasks remain unclear. This study conducts a comprehensive analysis of layer-wise visual representations in MLLMs, aiming to inform the selection of visual layers and guide the design of future visual fusion strategies.
3 Overall Setup
Problem Formulation
MLLMs typically comprise three core components: a vision encoder, a connector that maps visual features to the language space, and a large language model (LLM) Lin et al. (2024). This architecture empowers MLLMs to handle a diverse array of perception and reasoning tasks across both visual and textual modalities.
Most modern MLLMs adopt a pre-trained CLIP-ViT Radford et al. (2021) as their image encoder. A ViT encodes an image into a sequence of token embeddings through a stack of transformer blocks. Each block (or layer) progressively refines the visual representations, with earlier layers focusing on low-level spatial details and later layers capturing more abstract, semantic information.
Formally, given an image , the vision encoder produces a set of layer-wise outputs:
denotes the embedding at the -th layer, is the number of tokens, and is the dimension.
Despite the availability of rich multi-level features, most MLLMs select a single layer—often the penultimate or final one—to represent the entire image. This practice may overlook complementary signals from shallower layers that encode fine-grained visual details. In this work, we systematically investigate the impact of using different ViT layers for visual input and explore how selecting appropriate layers can improve MLLM performance across diverse tasks.
@@ 3 Overall Setup
Partitioning of Visual Representations
To examine the behavioral patterns of different visual layers, we analyze the relationships between them based on cosine similarities. Inspired by prior findings Sun et al. (2024) that LLMs exhibit several distinct representation spaces through such analysis, we similarly identify three significantly different representation spaces within CLIP-ViT.
As shown in Fig.1(a), three distinct representation spaces emerge among the visual layers (see Appendix A for computational details). Experiments on OCR and TextVQA (Fig. 1(b)) further indicate that shallow layers contribute little to performance, which rises substantially in the middle layers and peaks in the deep layers. Layers within the same representation space exhibit similar behaviors.
@@
Implementation Details
We employ CLIP ViT-L/14 (336px) Radford et al. (2021) as the visual encoder and 1.4B MobileLLaMA Chu et al. (2024) as the language model for efficiency analysis, with a one-layer MLP serving as the connector. Training follows a two-phase strategy aligned with LLaVA Liu et al. (2023b). AdamW optimizer with a cosine annealing scheduler is used, with learning rates of 1e-3 (phase one) and 2e-5 (phase two), and batch sizes of 256 and 128. Training on four NVIDIA A100 80GB GPUs takes 2 hours for phase one and 8 hours for phase two. We adopt the LLaVA 1.5 Liu et al. (2023b) dataset, comprising 558K image-caption pairs which is carefully filtered from CC3M Sharma et al. (2018) for pre-training and 665K conversational instances for instruction tuning. Unless explicitly noted, the experimental setup remains the same.
Evaluation Benchmarks
To comprehensively explore and evaluate various visual representations, we classified the benchmarks into four categories following previous work Tong et al. (2024a): General tasks, OCR tasks, Vision-centric tasks, and Hallucination tasks.
The General tasks category assesses basic vision-language reasoning abilities, including MME Fu et al. (2024) (yes/no questions on attributes like existence and color), MMBench Liu et al. (2024c) (multiple-choice across diverse aspects), SEEDBench Li et al. (2023a) (spatial and temporal reasoning), and GQA Hudson and Manning (2019a) (complex real-world VQA). The OCR category evaluates a model’s ability to recognize textual content from images, featuring TextVQA Singh et al. (2019) and OCRBench Liu et al. (2024d). The Vision-centric category emphasize fine-grained perception and localization, including CVBench Tong et al. (2024a) (evaluating spatial relations and depth), RealWorldQA (real-world QA), MMVet Yu et al. (2023) (general multimodal assessment), and RefCOCO Yu et al. (2016) (visual grounding). Finally, the Hallucination category includes POPE Li et al. (2023c), which evaluates whether MLLMs generate false or invented content not grounded in the image.
@@ 4 Experiment: Layer-wise Exploration
Previous studies have primarily used techniques such as linear probing and attention head decomposition to analyze CLIP-ViT representations Gandelsman et al. (2024). While these methods reveal what types of information are present in different ViT layers, they do not assess whether such information can be effectively utilized by MLLMs. The mere presence of information in a particular layer does not guarantee its usefulness when integrated into an MLLM. In contrast, our work goes beyond probing for representational content—we systematically evaluate how each ViT layer contributes to downstream MLLM performance. To this end, we conduct a layerwise exploration by individually connecting each visual layer to the language model, training the corresponding MLLM, and benchmarking its task performance. The layerwise performance is shown in Tab. 1 and Fig. 2.
4.1 Deep-to-Deep Layer Comparison
A common practice is to use deep layers from ViT as input to the MLLM. In this section, we investigate the effectiveness of this approach.
The final layer is not the optimal choice:
As shown in Tab. 1 and Fig. 2, the final layer does NOT perform the best on any benchmark. For general tasks, a noticeable performance drop is observed at the final layer, with OCR tasks exhibiting particularly severe degradation. A similar trend is evident in POPE. However, vision-centric tasks partially show this decline. Overall, these results indicate that the final layer is NOT the optimal choice for representation across tasks.